OpenStudy (akgorgeous):

ALGEBRA EQUATION- Given the system of equations presented here: 3x + 5y = 29 x + 4y = 16 Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated?

OpenStudy (akgorgeous):

Multiply the first equation by −1 to get −3x − 5y = −29 Multiply the second equation by −3 to get −3x − 12y = −48 Multiply the second equation by −1 to get −x − 4y = −16 Multiply the first equation by −3 to get −9x − 15y = −87

OpenStudy (sunnnystrong):

What do you think? If you have: \[3x + 5y = 29 \] \[x + 4y = 16\] & Your goal is to multiply one equation to create a system of equation where when combined --> a variable cancels.

OpenStudy (akgorgeous):

@3mar you got any ideas?

OpenStudy (3mar):

Where are you stuck?

OpenStudy (akgorgeous):

how to solve it.

OpenStudy (3mar):

We can use either substitution where we plug one equation into the other, or elimination where we combine the equations.

OpenStudy (mrs.ambrose614):

System of equations you have to eliminate 4y and5y

OpenStudy (3mar):

@akgorgeous Which way would you like to proceed with?

OpenStudy (akgorgeous):

elimination seems to be the easiest for me.

OpenStudy (3mar):

Nicer and easier choice! So with elimination you would to eliminate one variable from both equations, so you automatically got one equation with one variable!  How could you eliminate one variable?